Artificial rotating at an altitude of h = 300km above the Earth's surface
Find the throttle acceleration in the center of the constellation.
1
Expert's answer
2020-03-26T11:08:57-0400
"a=\frac{GM}{(R+h)^2}"

"a=\frac{(6.67\cdot10^{-11})(6\cdot10^{24})}{(6.4\cdot10^6+300\cdot10^3)^2}=8.9\frac{m}{s^2}"